# Emergent phases in bent-core patchy colloids

### v1

@@ I Introduction
In the present contribution we describe in detail the process that starts from the formation of BCs via the self-assembly of elliptic patchy colloids and ends up in the self-organization of these entities into BCLCs. Our design of the elliptic patchy particles envisages a binary mixture of mirror-symmetric units: while both species have the same elliptic shape, the two types of particles are mirror symmetric with respect to their patch decoration (leading thus to left- and right-handed particles). In fact, the two patches (per particle) are located close to the tip of the colloid. The form of the patch-patch interaction is based on the work of Russo et al. Russo et al. (2009), while the choice of the position, interaction range and interaction strength of the patches influences the features of the BC units. In a first step we demonstrate that our particles are indeed able to self-assemble via the related patch-patch bonding mechanism into stable BC entities; we demonstrate that our model is characterized by a high versatility, as the bonding angle can be tailored via the patch positions and/or the patch-patch interaction strength. The impact of all these parameters on the actual shape, stability and yield of these units is studied in detail. In a second step we consider an ensemble of such BCs and study their self-assembly into mesophases. Focusing for the moment on a particular set of model parameters we are able to trace out in a qualitative manner a diagram of states (in the temperature-density plane), which shows the emergence of nematic and smectic phases. All investigations have been carried out via extensive computer-simulations: we have applied both molecular dynamics (MD) and – in a complementary manner – Monte Carlo (MC) simulations, using different types of thermodynamic ensembles. As the proposed BC model can be easily extended to support BCs with different sets of features, it reveals itself as a powerful tool to gain a deeper insight into this fascinating class of functional materials with potential applications in optics, photonics, and sensing. The rest of our paper is organized as follows. Sec. II provides the theoretical framework which includes the model and the methodology to characterize the bent-cores. The computational details of the MD and MC procedures are provided in Sec. III. The results and discussion follow in Sec. IV. Finally, Sec. V provides the summary and conclusion.
@@ IV.2 Second step assembly stage
We now turn our attention to the ordering assemblies from the initial monomeric state. Figs. 5(a)-(d) show the time evolution of the morphology. The corresponding BC yields $f$ are shown in Fig. 5(e). At time $t=0$, when the system is quenched from $T=2$ to $1.2$, the system is in a homogeneous isotropic state where the monomers have not formed any BCs [Fig. 5(a)]. In the absence of any patches, this would remain an isotropic fluid, given our choice of the GB parameters. Due to the presence of attractive patches, opposite-handed monomers start adhering. The BC yield $f$ rapidly approaches 1 as time increases. An unusual observation from Figs. 5(c) and (d) is that the evolving ordered state has anti-ferroelectric smectic order SmZA Chen et al. (2023); Nacke et al. (2024). The latter has generated interest in the LC literature as a phase that fills an important void in the taxonomy of smectic LCs and has huge technological applications Lagerwall and Giesselmann (2006); Kim and Yoon (2018). We also evaluate the nematic and polar order parameters ($\mathcal{S}$ and $\mathcal{P}$) for the evolving morphology in Fig. 5(f). Note that the polar direction $\hat{p}$ of the neighbouring BCs in the ordered phase are antiparallel (or staggered). Consequently, we evaluate the polarization by considering only the magnitude of $\hat{p}$ along the direction of polarization ${\bf\boldsymbol{p}}$. In other words, we take the absolute value of $(\hat{p}_{i}\cdot\mathbf{p})$ before averaging over all BCs. As seen in Fig. 5(f), $\mathcal{S}$ and $\mathcal{P}$ are close to zero at early times as the system is in an isotropic state. As time evolves, there is the emergence of BCs and subsequent liquid crystalline order due to the GB interactions between the BC units. The BCs start forming small locally aligned regions, which then coalesce into larger domains with liquid crystalline order [Fig. 5(c) and (d)]. Consequently, there is an increase in the nematic and polar order parameters, as seen in Fig. 5(f). These observations, along with Figs. 5(e), emphasize the two-stage self-assembly in our model: the BC formation followed by long-range liquid crystalline order in the SmZA phase.
@@
The full set of BC states (phases) for a specific choice of model parameters ($\epsilon_{p}=15$, $\alpha=0.1$) in the optimal window is shown in the $\rho-T$ state diagram of Fig. 6(a). The different symbols indicate the different states observed in our simulations: isotropic state (indigo circles), antiferroelectric smectic SmZA (magenta triangles), arrested or frozen state (green squares) and disordered state (blue diamonds). The background color represents the BC yield. We see that there is a sizeable region in the $\rho-T$ space where the BC yield is greater than $0.9$. Panels (b) to (e) show typical morphologies obtained from MD (left) and MC (right) simulations for the four states obtained at indicated parameter values of the corresponding filled symbols in the state diagram. In (b), there is formation of BCs, but the system is in an isotropic state as the Gay-Berne attraction between them is not sufficient to overcome the thermal fluctuations. The region between the dashed lines is where we obtain the SmZA phase shown in the snapshot (c). For these parameter values, the BCs are fluid enough to self-assemble into large ordered domains. At low temperatures ($T<0.9$), there are locally aligned regions of BCs, but there is a lack of long-range order as these clusters are frozen. Such a kinetically arrested or frozen state due to low temperatures is shown in panel (d). At very high densities $\rho>0.6$, the particles lack the free volume to move and reorient, leading to a dense disordered BC state, as shown in panel (e). Similar morphologies are also observed in the NPT ensemble with pressure-tuned to achieve corresponding densities, see Fig. 7.
@@
In order to characterize the structure of the phases, we evaluate the directional pair correlation functions (PCF) resolved along axes in the local frame of reference of the BCs as described in section II.2.3. This is schematically illustrated in Fig. 8(a) where for each BC $i$, we compute the pair correlations with other BCs along parallel or perpendicular directions to its local alignment $\hat{n}_{i}$. Only those BCs that are within a finite region perpendicular to the direction of the correlation evaluation are included. The directional PCFs at $\rho=0.4$ and $T=1.5$, corresponding to the isotropic phase, are shown in Fig. 8(b). We observe only a few peaks in $g_{\perp}(r)$ and hardly any in $g_{\parallel}(r)$, indicating that at most, only two or three BCs stack along their short axis. In the case of the smectic anti-ferroelectric phase ($\rho=0.4$ and $T=1.2$), we observe multiple equidistant peaks along either direction, indicating the arrangement of BCs into multiple layers, as shown in Fig. 8(c). Interestingly, the correlations decay much more slowly in the direction of the long axis of the BCs ($g_{\perp}(r)$). In the frozen state at $\rho=0.4$ and $T=0.9$ [Fig. 8(e)], we find equidistant peaks that decay rapidly, corresponding to short-range correlations along either direction. This is due to the presence of multiple small domains of locally aligned BCs. In the dense disordered state ($\rho=0.7$ and $T=1.2$), we only find short-range correlations along the short axis of the BCs, as seen from Fig. 8(f). In all cases, our MD and MC results are in agreement with each other. The variation in the correlation lengths along the two directions ($l_{n}$ and $l_{p}$) with temperature $T$ is shown in Fig. 8(d), for two densities ($\rho=0.4$ and $0.7$). At the higher density, the system is in a dense disordered state, and hence the correlation lengths do not vary with temperature. At the lower density, we see a gradual increase in the correlation lengths with a peak centered around $T=1.2$, corresponding to the increased smectic anti-ferroelectric alignment, beyond which the correlations decay rapidly as the system enters the high-temperature isotropic phase. The plots also show that the size of the LC domains along the long axis of the BCs is larger than that in the orthogonal direction. The current investigation of parameters belonging to patch-patch interactions provides only isotropic and smectic anti-ferroelectric phases. This is because we have not fully explored the effect of patch positions on the formation of mesoscopic phases. Our preliminary explorations of the $\alpha$-$\theta_{1}$-$\theta_{2}$ space reveal a variety of phases, e.g., nematic and splay-bends, through appropriate choice of the model parameters.
@@ V Summary and Conclusion
We have developed a patchy ellipsoid model capable of generating mesoscopic liquid crystalline phases in a hierarchical two-stage self-assembly process, via the formation of BC dimers in the first stage. This is achieved through a binary mixture of ellipses having a pair of complementary and obliquely placed patches with specific combination rules to guide the formation BCs. The patches are modeled as points on the surface of the ellipses interacting via an inverted Gaussian potential. The second stage of the assembly into ordered liquid phases is governed by the Gay-Berne interactions acting between the ellipses themselves. Our model has three crucial parameters: (i) the patch interaction strength $\epsilon_{p}$, (ii) the patch interaction range $\alpha$ and (iii) the patch positions $\theta_{1}$ and $\theta_{2}$, that control the formation of the BCs, their mean bend angle and flexibility.
@@ Acknowledgements
Figure 6: (a) $\rho-T$ state diagram for the choice of parameters $\epsilon_{p}=15$, $\alpha=0.1$, in the NVT ensemble. The symbols correspond to different phases: isotropic (I, indigo circle), antiferroelectric smectic (SmZA, magenta triangle), arrested state due to low-temperature (AS1, green square), and arrested state due to high density (AS2, blue diamond). The background color represents the BC yield $f$ at the given state point. The region between dashed lines corresponds to SmZA. Snapshots of typical morphologies obtained from MD (left) and MC (right) simulations at state points marked with filled symbols in panel (a): (b) $T=1.5$, $\rho=0.4$ (I), (c) $T=1.2$, $\rho=0.4$ (SmZA), (d) $T=0.9$, $\rho=0.4$ (AS1), and (e) $T=1.2$, $\rho=0.7$ (AS2).
@@
Figure 8: (a) Schematic illustration of the evaluation of the directional PCF along the polar and nematic alignment of the central BC, with correlation lengths $l_{n}$ along nematic and $l_{p}$ along polar directions. Plots of the directional PCFs $g_{\parallel}(r_{\parallel})$ (in red), $g_{\perp}(r_{\perp})$ (in blue) vs. $r_{\parallel},r_{\perp}$ in the NVT ensemble for various values of $T$ and $\rho$: (b) $T=1.5$, $\rho=0.4$, (c) $T=1.2$, $\rho=0.4$, (e) $T=0.9$, $\rho=0.4$, and (f) $T=1.2$, $\rho=0.7$. All data correspond to the choice of model parameters: $\epsilon_{p}=15$, $\alpha=0.1$, $\theta_{1}=5$, and $\theta_{2}=10$. (d) Correlation lengths $l_{n}$ and $l_{p}$ as a function of temperature $T$ at densities $\rho$ = 0.4 and $\rho$ = 0.7.

### v2

# Two-stage assembly of patchy ellipses: From bent-core particles to liquid crystal analogs

@@ I Introduction
In the present contribution we describe in detail the process that starts from the formation of BCs via the self-assembly of elliptic patchy colloids and ends up in the self-organization of these entities into BCLCs. Our design of the elliptic patchy particles envisages a binary mixture of mirror-symmetric units: while both species have the same elliptic shape, the two types of particles are mirror symmetric with respect to their patch decoration (leading thus to left- and right-handed particles). In fact, the two patches (per particle) are located close to the tip of the colloid. The form of the patch-patch interaction is based on the work of Russo et al. Russo et al. (2009), while the choice of the position, interaction range and interaction strength of the patches influences the features of the BC units. In a first step we demonstrate that our particles are indeed able to self-assemble via the related patch-patch bonding mechanism into stable BC entities; we demonstrate that our model is characterized by a high versatility, as the bonding angle can be tailored via the patch positions and/or the patch-patch interaction strength. The impact of all these parameters on the actual shape, stability and yield of these units is studied in detail. In a second step we consider an ensemble of such BCs and study their self-assembly into mesophases. Focusing for the moment on a particular set of model parameters we are able to trace out in a qualitative manner a diagram of states (in the temperature-density plane), which shows the emergence of isotropic and smectic phases. All investigations have been carried out via extensive computer-simulations: we have applied both molecular dynamics (MD) and – in a complementary manner – Monte Carlo (MC) simulations, using different types of thermodynamic ensembles. As the proposed BC model can be easily extended to support BCs with different sets of features, it reveals itself as a powerful tool to gain a deeper insight into this fascinating class of functional materials with potential applications in optics, photonics, and sensing. The rest of our paper is organized as follows. Sec. II provides the theoretical framework which includes the model and the methodology to characterize the bent-cores. The computational details of the MD and MC procedures are provided in Sec. III. The results and discussion follow in Sec. IV. Finally, Sec. V provides the summary and conclusion.
@@ IV.2 Second step assembly stage
We now turn our attention to the ordering assemblies from the initial monomeric state. Figs. 5(a)-(d) show the time evolution of the morphology. The corresponding BC yields $f$ are shown in Fig. 5(e). At time $t=0$, when the system is quenched from $T=2$ to $1.2$, the system is in a homogeneous isotropic state where the monomers have not formed any BCs [Fig. 5(a)]. In the absence of any patches, this would remain an isotropic fluid, given our choice of the GB parameters. Due to the presence of attractive patches, opposite-handed monomers start adhering. The BC yield $f$ rapidly approaches 1 as time increases. An unusual observation from Figs. 5(c) and (d) is that the evolving ordered state has anti-ferroelectric smectic order. We also evaluate the nematic and polar order parameters ($\mathcal{S}$ and $\mathcal{P}$) for the evolving morphology in Fig. 5(f). Note that the polar direction $\hat{p}$ of the neighbouring BCs in the ordered phase are antiparallel (or staggered). Consequently, we evaluate the polarization by considering only the magnitude of $\hat{p}$ along the direction of polarization ${\bf\boldsymbol{p}}$. In other words, we take the absolute value of $(\hat{p}_{i}\cdot\mathbf{p})$ before averaging over all BCs. As seen in Fig. 5(f), $\mathcal{S}$ is close to zero at early times as the system is in an isotropic state. As time evolves, there is the emergence of BCs and subsequent liquid crystalline order due to the GB interactions between the BC units. The BCs start forming small locally aligned regions, which then coalesce into larger domains with liquid crystalline order [Fig. 5(c) and (d)]. Consequently, there is an increase in the nematic and polar order parameters, as seen in Fig. 5(f). These observations, along with Figs. 5(e), emphasize the two-stage self-assembly in our model: the BC formation followed by long-range liquid crystalline order in the smectic anti-ferroelectric phase.
@@
The full set of BC states (phases) for a specific choice of model parameters ($\epsilon_{p}=15$, $\alpha=0.1$) in the optimal window is shown in the $\rho-T$ state diagram of Fig. 6(a). The different symbols indicate the different states observed in our simulations: isotropic state (indigo circles), anti-ferroelectric smectic (magenta triangles), arrested or frozen state (green squares) and disordered state (blue diamonds). The background color represents the BC yield. We see that there is a sizeable region in the $\rho-T$ space where the BC yield is greater than $0.9$. Panels (b) to (e) show typical morphologies obtained from MD (left) and MC (right) simulations for the four states obtained at indicated parameter values of the corresponding filled symbols in the state diagram. In (b), there is formation of BCs, but the system is in an isotropic state as the Gay-Berne attraction between them is not sufficient to overcome the thermal fluctuations. The region between the dashed lines is where we obtain the smectic anti-ferroelectric phase shown in the snapshot (c). For these parameter values, the BCs are fluid enough to self-assemble into large ordered domains. At low temperatures ($T<0.9$), there are locally aligned regions of BCs, but there is a lack of long-range order as these clusters are frozen. Such a kinetically arrested or frozen state due to low temperatures is shown in panel (d). At very high densities $\rho>0.6$, the particles lack the free volume to move and reorient, leading to a dense disordered BC state, as shown in panel (e). Similar morphologies are also observed in the NPT ensemble with pressure-tuned to achieve corresponding densities, see Fig. 7.
@@
In order to characterize the structure of the phases, we evaluate the directional radial distribution functions (RDF) resolved along axes in the local frame of reference of the BCs as described in section II.2.3. This is schematically illustrated in Fig. 8(a) where for each BC $i$, we compute the RDF with other BCs along parallel or perpendicular directions to its local alignment $\hat{n}_{i}$. Only those BCs that are within a finite region perpendicular to the direction of the correlation evaluation are included. The directional RDFs at $\rho=0.4$ and $T=1.5$, corresponding to the isotropic phase, are shown in Fig. 8(b). We observe only a few peaks in $g_{\perp}(r_{\perp})$ and hardly any in $g_{\parallel}(r_{\parallel})$, indicating that at most, only two or three BCs stack along their short axis. In the case of the smectic anti-ferroelectric phase ($\rho=0.4$ and $T=1.2$), we observe multiple equidistant peaks along either direction, indicating the arrangement of BCs into multiple layers, as shown in Fig. 8(c). Interestingly, the correlations decay much more slowly in the direction of the long axis of the BCs ($g_{\perp}(r_{\perp})$). In the frozen state at $\rho=0.4$ and $T=0.9$ [Fig. 8(e)], we find equidistant peaks that decay rapidly, corresponding to short-range correlations along either direction. This is due to the presence of multiple small domains of locally aligned BCs. In the dense disordered state ($\rho=0.7$ and $T=1.2$), we only find short-range correlations along the short axis of the BCs, as seen from Fig. 8(f). In all cases, our MD and MC results are in agreement with each other. The variation in the correlation lengths along the two directions ($l_{n}$ and $l_{p}$) with temperature $T$ is shown in Fig. 8(d), for two densities ($\rho=0.4$ and $0.7$). At the higher density, the system is in a dense disordered state, and hence the correlation lengths do not vary with temperature. At the lower density, we see a gradual increase in the correlation lengths with a peak centered around $T=1.2$, corresponding to the increased smectic anti-ferroelectric alignment, beyond which the correlations decay rapidly as the system enters the high-temperature isotropic phase. The plots also show that the size of the LC domains along the long axis of the BCs is larger than that in the orthogonal direction. The current investigation of parameters belonging to patch-patch interactions provides only isotropic and smectic anti-ferroelectric phases. This is because we have not fully explored the effect of patch positions on the formation of mesoscopic phases. Our preliminary explorations of the $\alpha$-$\theta_{1}$-$\theta_{2}$ space reveal a variety of phases, e.g., nematic and splay-bends, through appropriate choice of the model parameters.
@@ V Summary and Conclusion
We have developed a patchy ellipsoid model capable of generating mesoscopic liquid crystalline phases in a hierarchical two-stage self-assembly process, via the formation of BC dimers in the first stage. This is achieved through a binary mixture of ellipses having a pair of complementary and obliquely placed patches with specific combination rules to guide the formation of BCs. The patches are modeled as points on the surface of the ellipses interacting via an inverted Gaussian potential. The second stage of the assembly into ordered liquid phases is governed by the Gay-Berne interactions acting between the ellipses themselves. Our model has three crucial parameters: (i) the patch interaction strength $\epsilon_{p}$, (ii) the patch interaction range $\alpha$ and (iii) the patch positions $\theta_{1}$ and $\theta_{2}$, that control the formation of the BCs, their mean bend angle and flexibility.
@@ Acknowledgements
Figure 6: (a) $\rho-T$ state diagram for the choice of parameters $\epsilon_{p}=15$, $\alpha=0.1$, in the NVT ensemble. The symbols correspond to different phases: isotropic (I, indigo circle), anti-ferroelectric smectic (magenta triangle), arrested state due to low-temperature (AS1, green square), and arrested state due to high density (AS2, blue diamond). The background color represents the BC yield $f$ at the given state point. The region between dashed lines corresponds to smectic anti-ferroelectric. Snapshots of typical morphologies obtained from MD (left) and MC (right) simulations at state points marked with filled symbols in panel (a): (b) $T=1.5$, $\rho=0.4$ (I), (c) $T=1.2$, $\rho=0.4$ (smectic anti-ferroelectric), (d) $T=0.9$, $\rho=0.4$ (AS1), and (e) $T=1.2$, $\rho=0.7$ (AS2).
@@
Figure 8: (a) Schematic illustration of the evaluation of the directional RDFs along the polar and nematic alignment of the central BC, with correlation lengths $l_{n}$ along nematic and $l_{p}$ along polar directions. Plots of the directional RDFs $g_{\parallel}(r_{\parallel})$ (in red), $g_{\perp}(r_{\perp})$ (in blue) vs. $r_{\parallel},r_{\perp}$ in the NVT ensemble for various values of $T$ and $\rho$: (b) $T=1.5$, $\rho=0.4$, (c) $T=1.2$, $\rho=0.4$, (e) $T=0.9$, $\rho=0.4$, and (f) $T=1.2$, $\rho=0.7$. All data correspond to the choice of model parameters: $\epsilon_{p}=15$, $\alpha=0.1$, $\theta_{1}=5$, and $\theta_{2}=10$. (d) Correlation lengths $l_{n}$ and $l_{p}$ as a function of temperature $T$ at densities $\rho$ = 0.4 and $\rho$ = 0.7.